OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

hartnn (hartnn):

divide both sides by -3 what do u get ?

OpenStudy (jdoe0001):

\(\bf -3|x + 7| = -12\implies |x + 7| =\cfrac{-12}{-3} \\ \quad \\ \implies \begin{cases} +(x+7)=\cfrac{-12}{-3} \\ \quad \\ \bf -(x+7)=\cfrac{-12}{-3} \end{cases}\) solve each scenario/case for "x", thus 2 values for "x", one for each case

OpenStudy (anonymous):

x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

hartnn (hartnn):

whats -12/-3 = ... ?

OpenStudy (anonymous):

4 ?

hartnn (hartnn):

yes! when |a| = b , we get a=b or a =-b so when |x+7| = 4 what do u get ?

OpenStudy (anonymous):

Uhm....- 3 ?

hartnn (hartnn):

thats what you got for x+7 =4 x =-3 is correct :) also find x from x+7 = -4 subtract 7 from both sides, what u get ?

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

so it would be x = −3, x = −11

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

Thank you I have like three more questions on my pre quiz, its not even being graded but could you still help out with them ?

hartnn (hartnn):

sure :) just make sure to ask new question on new post, so that if i am unavailable, others will help you :)

OpenStudy (anonymous):

Okay! I'll close this question.